Question 8 What is the oxidation state (charge) on Mn in LiMnO2 ? Answer Enter your answer here
Answer #1

The oxidation state of Mn in LiMnO2 =

Explanation:

LiMnO2 :

+1 + X + 2(-2) = 0

+1+X-4 = 0

X-3 = 0

X = +3

Therefore, the oxidation state of Mn in LiMnO2 is +3.